THE A^{6}Σ^{+}− X^{6}Σ^{+} TRANSITION OF CrH, EINSTEIN COEFFICIENTS AND AN IMPROVED DESCRIP-TION OF THE A STATE

C. W. BAUSCHLICHER, JR. , Space Technology Division, NASA Ames Research Center, Moffett Field, CA 94035; \underline{R. S. RAM}, Department of Chemistry, University of Arizona, Tucson, AZ 85721; P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1; C. G. PARSONS and D. GALEHOUSE, Department of Physics, University of Dayton, Dayton, OH 45469.